# Multimodal Data Integration for Oncology in the Era of Deep Neural Networks: A Review

### v2

@@ 3.3.3 Molecular Data
Graphs are a natural choice for representing molecular data such as omic-centric (DNA, RNA, or proteins) or single-cell centric. Individual modalities are processed separately to generate graph representations that are then processed through GNNs followed by the classifier to predict the downstream task, as illustrated in Fig. 8. One method of representing proteins as graphs is to depict the amino acid residue in the protein as the node and the relationship between residues denoted by edge [87]. The residue information is depicted as node embedding, whereas the relational information between two residues is represented as the edge feature vector. Fout et al. [87] used spatial ConvGNNs to predict interfaces between proteins which is important in drug discovery problems. Deep predictor of drug-drug interactions (DPDDI) predicted the drug-drug interactions using GCN followed by a 5-layer classical neural network [88]. Molecular pre-training graph net (MPG) [89] is a powerful framework based on GNN and BERT to learn drug-drug and drug-target interactions. Graph-based Attention Model (GRAM) [90] handled the data inefficiency by supplementing EHRs with hierarchical knowledge in the medical ontology. A few recent works have applied GNNs to single-cell data. scGCN [91] is a knowledge transfer framework in single-cell omics data such as mRNA or DNA. scGNN [92] processed cell-cell relations through GNNs for the task of missing-data imputation and cell clustering on single-cell RNA sequencing (scRNA-seq) data.
3.4 Multimodal Pre-learning Fusion
@@
The first and most primitive form of MML is the pre-learning fusion (see Fig. 4), where features extracted from individual modalities of data are merged, and the fused representations are then used for training the multimodal primary learner model. In the context of GNNs being the primary learning model, the extraction step of individual modality representations can be hand-engineered (e.g., dimensionality reduction) or learned by DL models (e.g., CNNs, Transformers). Cui et al. [93] proposed a GNN-based early fusion framework to learn latent representations from radiological and clinical modalities for Lymph node metastasis (LNM) prediction in esophageal squamous cell carcinoma (ESCC). The extracted features from the two modalities using UNet and CNN-based encoders were fused together with category-wise attention as node representation. The message passing from conventional GAT and correlation-based GAT learned the neighborhood weights. The attention attributes were used to update the final node features before classification by a 3-layer fully connected network. For Autism spectrum disorder, Alzheimer’s disease, and ocular diseases, a multimodal learning framework called Edge-Variational GCN (EV-GCN) [94] fuses the radiology features extracted from fMRI images with clinical feature vectors for each patient. An MLP-based pairwise association encoder is used to fuse the input feature vectors and to generate the edge weights of the population graph. The partially labeled population graph is then processed through GCN layers to generate the diagnostic graph of patients.
@@
Figure 8: Graph processing pipeline on non-imagery data. Adapted from [95].
3.5 Multimodal Cross-learning Fusion
@@
Cross-MML involves intermediate fusion and/or cross-links among the models being trained on individual modalities (see Fig. 4). For this survey, we consider the GNN-based hierarchical learning mechanisms as the cross-MML methods. Hierarchical frameworks involve learning for one modality and using the learned latent embeddings in tandem with other data modalities sequentially to get the final desired low-dimensional representations. Lian et al. [96] used a sequential learning framework where tumor features learned from CT images using the ViT model were used as node features of the patient population graph for subsequent processing by the GraphSAGE model. The hierarchical learning from radiological and clinical data using Transformer-GNN outperformed the ResNet-Graph framework in survival prediction of early-stage NSCLC. scMoGNN [97] is the first method to apply GNNs in multimodal single-cell data integration using a cross-learning fusion-based GNN framework. Officially winning first place in modality prediction task at the NeurIPS 2021 competition, scMoGNN showed superior performance on various tasks by using paired data to generate cell-feature graphs. Hierarchical cell-to-tissue-graph network (HACT-Net) combined the low-level cell-graph features with the high-level tissue-graph features through two hierarchical GINs on breast cancer multi-class prediction [98]. Data imputation, a method of populating the missing values or false zero counts in single-cell data mostly done using DL autoencoders (AE) architecture, has recently been accomplished using GNNs. scGNN [92] used imputation AE and graph AE in an iterative manner for imputation, and GraphSCI [99] used GCN with AE to impute the single-cell RNA-seq data using the cross-learning fusion between the GCN and the AE networks. Clustering is a method of characterizing cell types within a tissue sample. Graph-SCC [100] clustered cells based on scRNA-seq data through self-supervised cross-learning between GCN and a denoising AE network. Recently a multilayer GNN framework, Explainable Multilayer GNN (EMGNN), has been proposed for cancer gene prediction task using multi-omics data from 16 different cancer types [101].
3.6 Multimodal Post-learning
@@
Because of their superior performance, the most sought-after multimodal fusion technique in the literature is the post-learning fusion method, where the individual data views are first processed by the primary learner and later fused for the downstream predictive task [102]. In the post-learning fusion paradigm, the hand-crafted features perform better than the deep features when the dimensionality of input data is low, and vice versa [102]. Many interesting GNN-based works involving the post-learning fusion mechanism have recently been published. Decagon [103] used a multimodal approach on GCNs using proteins and drugs interactions to predict exact side effects as a multi-relational link prediction task. Drug–target affinity (DTA) [104] experimented with four different flavors of GNNs (GCN, GAT, GIN, GAT-GCN) along with a CNN to fuse together molecular embeddings and proteins sequences for predicting drug-target affinity. PathomicFusion [75] combined the morphological features extracted from image patches (using CNNs), cell-graph features from cell-graphs of histology images (GraphSAGE-based GCNs), and genomic features (using feed-forward network) for survival prediction on glioma and clear cell renal cell carcinoma. Shi et al. [105] proposed a late-fusion technique to study screening of cervical cancer at early stages, by using CNNs to extract features from histology images, followed by K-means clustering to generate graphs which are processed through two-layer GCN. BDR-CNN-GCN (batch normalized, dropout, rank-based pooling) [106] used the same mammographic images to extract image-level features using CNN and relation-aware features using GCN. The two feature sets are fused using dot product followed by a trainable linear projection for the breast cancer classification. Under the umbrella of multi-omics data, many GNN-based frameworks have been proposed recently. Molecular omics network(MOOMIN)[107], a multi-modal heterogeneous GNN to predict oncology drug combinations, processed molecular structure, protein features, and cell lines through GCN-based encoders, followed by late-fusion using a bipartite drug-protein interaction graph. Multi-omics graph convolutional networks (MOGONET) [95] used a GCN-GAN late fusion technique for the classification of four different diseases, including three cancer types, breast, kidney, and glioma. Leng et al. [108] extended MOGONET to benchmark three multi-omics datasets on two different tasks using sixteen DL networks and concluded that GAT-based GNN had the best classification performance. Multi-Omics Graph Contrastive Learner(MOGCL) [109] used graph structure and contrastive learning information to generate representations for improved downstream classification tasks on the breast cancer multi-omics dataset using late-fusion. Similar to MOGCL, Park et al. [110] developed a GNN-based multi-omics model that integrated mRNA expression, DNA methylation, and DNA sequencing data for NSCLC diagnosis.
4 Transformers in Multi-modal Learning
@@
Transformers can be viewed as a type of GNN [10]. Transformers are attention-based DNN models that originally gained attention in NLP [111]. Transformers implement scaled dot-product of the input with itself and can process various types of data in parallel [111]. Transformers can handle sequential data and learn long-range dependencies, making them well-suited for tasks such as language translation, language modeling, question answering, and many more [112]. Unlike Recurrent Neural Networks (RNNs) and CNNs, Transformers use self-attention operations to weigh the importance of different input tokens (or embeddings) at each time step. This allows them to handle sequences of arbitrary length and to capture dependencies between input tokens that are far apart in the sequence [111]. Transformer models have also been applied to other data types, such as images [113], audio [114], and time-series analysis [115], resulting in a new wave of multi-modal applications. Transformers can handle input sequences of different modalities in a unified way, using the same self-attention mechanism which processes the inputs as a fully connected graph [10]. This allows Transformers to capture complex dependencies between different modalities, such as visual and textual information in visual question-answering (VQA) tasks [116]. Pre-training Transformers on large amounts of data, using unsupervised or self-supervised learning, and then fine-tuning for specific downstream tasks, has led to the development of foundation models, such as BERT [33], GPT [117], RoBERTa [34], CLIP [5], T5 [118], BART [119], BLOOM [120], ALIGN [121], CoCa [122] and more.
@@
Multimodal Transformers are a recent development in the field of MML, which extends the capabilities of traditional Transformers to handle multiple data modalities. The inter-modality dependencies are captured by the cross-attention mechanism in multimodal Transformers, allowing the model to jointly reason and extract rich data representations. There are various types of multimodal Transformers, such as Unified Transformer (UniT) [123], Multi-way Multimodal Transformer (MMT) [124], CLIP [5], Flamingo [125], CoCa [122], Perceiver IO [126], and GPT-4[6].
Figure 9: The original Transformer architecture [111]. A Transformer has multiple encoders and decoders blocks based on application and data type.
4.1 The Transformer Architecture
@@
The original Transformer (Fig. 9) was composed of encoder/ decoder blocks, each made up of several layers of self-attention and feed-forward neural networks. The encoder takes the input sequence and generates hidden representations, which are then fed to the decoder. The decoder generates the output sequence by attending to the encoder’s hidden representations and the previous tokens (i.e., auto-regressive). The self-attention operation (or scaled dot-product) is a crucial component of Transformer. It determines the significance of each element in the input sequence with respect to the whole input. Self-attention operates by computing a weighted sum of the input sequence’s hidden representations, where the weights are determined by the dot product between the query vector and the key vector, followed by a scaling operation to stabilize the gradients. The resulting weighted sum is multiplied by a value vector to obtain the output of the self-attention operation. There has been a tremendous amount of work on various facets of Transformer architecture. The readers are referred to relevant review papers [112, 10, 127, 128].
@@ 4.2 Multimodal Transformers
The self-attention allows a Transformer model to process each input as a fully connected graph and attend to (or equivalently learn from) the global patterns present in the input. This makes Transformers compatible with various data modalities by treating each token (or its embedding) as a node in the graph. To use Transformers for a data modality, we need to tokenize the input and select an embedding space for the tokens. Tokenization and embedding selections are flexible and can be done at multiple granularity levels, such as using raw features, ML-extracted features, patches from the input image, or graph nodes. Table 1 summarizes some common practices related to these aspects in oncology.
@@
Modalities Tokenization Token Embeddings
Pathology Patch CNNs [129]
Radiology Patch CNNs [130]
Electronic Health Record ICD Code
@@
Handling inter-modality interactions is the main challenge in developing multimodal Transformer models. Usually, it is done through one of these fusion methods; early fusion of data modalities, cross-attention, hierarchical attention, and late fusion, as illustrated in Fig. 10. In the following, we present and compare data processing steps for these four methods using two data modalities as an example. The same analysis can be extended to multiple modalities.
@@ 4.2.1 Early Fusion
Early fusion is the simplest way to combine data from multiple modalities. The data from different modalities are concatenated to a single input before feeding it to the Transformer model which processes input as a single entity. Mathematically, the concatenation operation is: $x_{cat}$=$[x_{1},x_{2}]$, where $x_{1}$ and $x_{2}$ are the inputs from two data modalities, and $x_{cat}$ is the concatenated input to the model. Early fusion is simple and efficient, however, it assumes that all modalities are equally important and relevant for the task at hand [136], which may not always be practically true [137].
@@ 4.2.2 Cross-Attention Fusion
Cross-attention is a relatively more flexible approach to model the interactions between data modalities and learn their joint representations. The Transformer layers attend to different modalities at different stages of data processing. Cross-attention allows the model to selectively attend to different modalities based on their relevance to the task [138], and capture complex interactions between the modalities [139].
@@ 4.2.3 Hierarchical Fusion
Hierarchical fusion is a complex approach to combine multiple modalities. For instance, DFTR (Depth-supervised Fusion Transformer for Salient Object Detection) employs hierarchical feature extraction to improve salient object detection performance by fusing low-level spatial features and high-level semantic features from different scales [140]. Yang et al.[141], introduced a hierarchical approach to fine-grained classification using a fusion transformer. Furthermore, the Hierarchical Multimodal Transformer (HMT) for video summarization can capture global dependencies and multi-hop relationships among video frames using a transformer model [142].
@@ 4.2.4 Late Fusion
In late fusion, each data modality is processed independently by its Transformer model, and the branch outputs are concatenated and passed through the final classifier. Late fusion allows the model to capture the unique features of each modality while still learning their joint representation. Sun et al. [143] proposed a multi-modal adaptive late fusion transformer network for estimating the levels of depression. Their model independently extracts long-term temporal information from audio and visual data and then fuses weights at the end to learn joint representation.
Figure 10: Four different strategies of fusing information from various data modalities in multimodal Transformers.
4.3 Transformers for Oncology Data Processing
Transformers have been successfully applied to various tasks in oncology, including cancer screening, diagnosis, prognosis, treatment selection, and prediction of clinical variables [9, 144, 145, 96, 129]. For instance, a Transformer-based model was used to predict the presence and grade of breast cancer using a combination of imaging and genomics data [9]. Transmil [144] was a transformer model proposed to process histopathology images using self-attention to learn and classify the correlation between multiple instances. Recently, a Transformer and convolution parallel network, TransConv [145], was proposed for automatic brain tumor segmentation using MRI data. Transformers and GNNs have also been combined in MML for early-stage NSCLC prognostic prediction using the patient’s clinical and pathological features and by modeling the patient’s physiological network [96]. Similarly, a multimodal co-attention Transformer was proposed for survival prediction using WSIs[129] by learning from the image and genomic sequences and using the co-attention mechanism to learn the interactions between the two data modalities [129].
5 Challenges in Multimodal Learning (MML)

### v3

@@ 2.3.3 Molecular Datasets
Graphs are a natural choice for representing molecular data such as omic-centric (DNA, RNA, or proteins) or single-cell centric. Individual modalities are processed separately to generate graph representations that are then processed through GNNs followed by the classifier to predict the downstream task, as illustrated in Figure 7. One method of representing proteins as graphs is to depict the amino acid residue in the protein as the node and the relationship between residues denoted by edge [101]. The residue information is depicted as node embedding, whereas the relational information between two residues is represented as the edge feature vector. Fout et al. [101] used spatial ConvGNNs to predict interfaces between proteins which is important in drug discovery problems. Deep predictor of drug-drug interactions (DPDDI) predicted the drug-drug interactions using GCN followed by a 5-layer classical neural network [102]. Molecular pre-training graph net (MPG) [103] is a powerful framework based on GNN and Bidirectional Encoder Representations from Transformers (BERT) to learn drug-drug and drug-target interactions. Graph-based Attention Model (GRAM) [104] handled the data inefficiency by supplementing EHRs with hierarchical knowledge in the medical ontology. A few recent works have applied GNNs to single-cell data. scGCN [105] is a knowledge transfer framework in single-cell omics data such as mRNA or DNA. scGNN [106] processed cell-cell relations through GNNs for the task of missing-data imputation and cell clustering on single-cell RNA sequencing (scRNA-seq) data.
2.4 MML - Data Fusion at the Pre-Learning Stage
@@
The first and most primitive form of MML is the pre-learning fusion (see Figure 3), where features extracted from individual modalities of data are merged, and the fused representations are then used for training the multimodal primary learner model. In the context of GNNs being the primary learning model, the extraction step of individual modality representations can be hand-engineered (e.g., dimensionality reduction) or learned by DL models (e.g., CNNs, Transformers). Cui et al. [107] proposed a GNN-based early fusion framework to learn latent representations from radiological and clinical modalities for Lymph node metastasis (LNM) prediction in esophageal squamous cell carcinoma (ESCC). The extracted features from the two modalities using UNet and CNN-based encoders were fused together with category-wise attention as node representation. The message passing from conventional GAT and correlation-based GAT learned the neighborhood weights. The attention attributes were used to update the final node features before classification by a 3-layer fully connected network. For Autism spectrum disorder, Alzheimer’s disease, and ocular diseases, a multimodal learning framework called Edge-Variational GCN (EV-GCN) [108] fuses the radiology features extracted from fMRI images with clinical feature vectors for each patient. An MLP-based pairwise association encoder is used to fuse the input feature vectors and to generate the edge weights of the population graph. The partially labeled population graph is then processed through GCN layers to generate the diagnostic graph of patients.
@@
Figure 7: Graph data processing pipeline for non-imagery data, including molecular and textual data. Adapted from [109]. Abbreviations used: GNN - graph neural network, FC - Fully-Connected, MLP - Multi-Layer Perception.
2.5 MML - Data Fusion using Cross-Modality Learning
@@
Cross-MML involves intermediate fusion and/or cross-links among the models being trained on individual modalities (see Figure 3). For this survey, we consider the GNN-based hierarchical learning mechanisms as the cross-MML methods. Hierarchical frameworks involve learning for one modality and using the learned latent embeddings in tandem with other data modalities sequentially to get the final desired low-dimensional representations. Lian et al. [110] used a sequential learning framework where tumor features learned from CT images using the ViT model were used as node features of the patient population graph for subsequent processing by the GraphSAGE model. The hierarchical learning from radiological and clinical data using Transformer-GNN outperformed the ResNet-Graph framework in survival prediction of early-stage NSCLC. scMoGNN [111] is the first method to apply GNNs in multimodal single-cell data integration using a cross-learning fusion-based GNN framework. Officially winning first place in modality prediction task at the NeurIPS 2021 competition, scMoGNN showed superior performance on various tasks by using paired data to generate cell-feature graphs. Hierarchical cell-to-tissue-graph network (HACT-Net) combined the low-level cell-graph features with the high-level tissue-graph features through two hierarchical GINs on breast cancer multi-class prediction [112]. Data imputation, a method of populating the missing values or false zero counts in single-cell data mostly done using DL autoencoders (AE) architecture, has recently been accomplished using GNNs. scGNN [106] used imputation AE and graph AE in an iterative manner for imputation, and GraphSCI [113] used GCN with AE to impute the single-cell RNA-seq data using the cross-learning fusion between the GCN and the AE networks. Clustering is a method of characterizing cell types within a tissue sample. Graph-SCC [114] clustered cells based on scRNA-seq data through self-supervised cross-learning between GCN and a denoising AE network. Recently, a multilayer GNN framework, Explainable Multilayer GNN (EMGNN), has been proposed for cancer gene prediction tasks using multi-omics data from 16 different cancer types [115].
2.6 MML - Data Fusion in Post-Learning Regime
@@
Post-learning fusion methods include processing individual data modalities and later fusing them for the downstream predictive task [116]. In the post-learning fusion paradigm, the hand-crafted features perform better than the deep features when the dimensionality of input data is low, and vice versa [116]. Many interesting GNN-based works involving the post-learning fusion mechanism have recently been published. Decagon [117] used a multimodal approach on GCNs using proteins and drug interactions to predict exact side effects as a multi-relational link prediction task. Drug–target affinity (DTA) [118] experimented with four different flavors of GNNs (GCN, GAT, GIN, GAT-GCN) along with a CNN to fuse together molecular embeddings and protein sequences for predicting drug-target affinity. PathomicFusion [90] combined the morphological features extracted from image patches (using CNNs), cell-graph features from cell-graphs of histology images (GraphSAGE-based GCNs), and genomic features (using a feed-forward network) for survival prediction on glioma and clear cell renal cell carcinoma. Shi et al. [119] proposed a late-fusion technique to study screening of cervical cancer at early stages by using CNNs to extract features from histology images, followed by K-means clustering to generate graphs which are processed through two-layer GCN. BDR-CNN-GCN (batch normalized, dropout, rank-based pooling) [120] used the same mammographic images to extract image-level features using CNN and relation-aware features using GCN. The two feature sets are fused using a dot product followed by a trainable linear projection for breast cancer classification. Under the umbrella of multi-omics data, many GNN-based frameworks have been proposed recently. Molecular omics network(MOOMIN)[121], a multi-modal heterogeneous GNN to predict oncology drug combinations, processed molecular structure, protein features, and cell lines through GCN-based encoders, followed by late-fusion using a bipartite drug-protein interaction graph. Multi-omics graph convolutional networks (MOGONET) [109] used a GCN-GAN late fusion technique for the classification of four different diseases, including three cancer types: breast, kidney, and glioma. Leng et al. [122] extended MOGONET to benchmark three multi-omics datasets on two different tasks using sixteen DL networks and concluded that GAT-based GNN had the best classification performance. Multi-Omics Graph Contrastive Learner(MOGCL) [123] used graph structure and contrastive learning information to generate representations for improved downstream classification tasks on the breast cancer multi-omics dataset using late-fusion. Similar to MOGCL, Park et al. [124] developed a GNN-based multi-omics model that integrated mRNA expression, DNA methylation, and DNA sequencing data for NSCLC diagnosis.
3 Transformers in MML
@@
Transformers are attention-based DNN models originally proposed for NLP [125]. Transformers implement scaled dot-product of the input with itself and can process various types of data in parallel [125]. Transformers can handle sequential data and learn long-range dependencies, making them well-suited for tasks such as language translation, language modeling, question answering, and many more [126]. Unlike Recurrent Neural Networks (RNNs) and CNNs, Transformers use self-attention operations to weigh the importance of different input tokens (or embeddings) at each time step. This allows them to handle sequences of arbitrary length and to capture dependencies between input tokens that are far apart in the sequence [125]. Transformers can be viewed as a type of GNN [30]. Transformers are used to process other data types, such as images [127], audio [128], and time-series analysis [129], resulting in a new wave of multi-modal applications. Transformers can handle input sequences of different modalities in a unified way, using the same self-attention mechanism, which processes the inputs as a fully connected graph [30]. This allows Transformers to capture complex dependencies between different modalities, such as visual and textual information in visual question-answering (VQA) tasks [130].
@@
Pre-training Transformers on large amounts of data, using unsupervised or self-supervised learning, and then fine-tuning for specific downstream tasks, has led to the development of foundation models [29], such as BERT [50], GPT [131], RoBERTa [51], CLIP [23], T5 [132], BART [133], BLOOM [134], ALIGN [135], CoCa [136] and more. Multimodal Transformers are a recent development in the field of MML, which extends the capabilities of traditional Transformers to handle multiple data modalities. The inter-modality dependencies are captured by the cross-attention mechanism in multimodal Transformers, allowing the model to jointly reason and extract rich data representations. There are various types of multimodal Transformers, such as Unified Transformer (UniT) [137], Multi-way Multimodal Transformer (MMT) [138], CLIP [23], Flamingo [139], CoCa [136], Perceiver IO [140], and GPT-4[21].
Figure 8: The original Transformer architecture is presented [125]. A Transformer can have multiple encoder and decoder blocks, as well as some additional layers.
3.1 Model Architecture
@@
The original Transformer (Figure 8) was composed of multiple encoder and decoder blocks, each made up of several layers of self-attention and feed-forward neural networks. The encoder takes the input sequence and generates hidden representations, which are then fed to the decoder. The decoder generates the output sequence by attending to the encoder’s hidden representations and the previous tokens (i.e., auto-regressive). The self-attention operation (or scaled dot-product) is a crucial component of the Transformer. It determines the significance of each element in the input sequence with respect to the whole input. Self-attention operates by computing a weighted sum of the input sequence’s hidden representations, where the weights are determined by the dot product between the query vector and the key vector, followed by a scaling operation to stabilize the gradients. The resulting weighted sum is multiplied by a value vector to obtain the output of the self-attention operation. There has been a tremendous amount of work on various facets of Transformer architecture. The readers are referred to relevant review papers [126, 30, 141, 142].
@@ 3.2 Multimodal Transformers
Self-attention allows a Transformer model to process each input as a fully connected graph and attend to (or equivalently learn from) the global patterns present in the input. This makes Transformers compatible with various data modalities by treating each token (or its embedding) as a node in the graph. To use Transformers for a data modality, we need to tokenize the input and select an embedding space for the tokens. Tokenization and embedding selections are flexible and can be done at multiple granularity levels, such as using raw features, ML-extracted features, patches from the input image, or graph nodes. Table 1 summarizes some common practices used for various data types in oncology datasets. Handling inter-modality interactions is the main challenge in developing multimodal Transformer models. Usually, it is done through one of these fusion methods: early fusion of data modalities, cross-attention, hierarchical attention, and late fusion, as illustrated in Figure 9. In the following, we present and compare data processing steps for these four methods using two data modalities as an example. The same analysis can be extended to multiple modalities.
@@
Data Modalities Tokenization Level Token Embeddings Model
Pathology images Patch CNNs [143]
Radiology images Patch CNNs [144]
EHR data ICD code
@@ 3.2.1 Early Fusion
Early fusion is the simplest way to combine data from multiple modalities. The data from different modalities are concatenated to a single input before being fed to the Transformer model, which processes the input as a single entity. Mathematically, the concatenation operation is represented as $x_{cat}$=$[x_{1},x_{2}]$, where $x_{1}$ and $x_{2}$ are the inputs from two data modalities, and $x_{cat}$ is the concatenated input to the model. Early fusion is simple and efficient. However, it assumes that all modalities are equally important and relevant for the task at hand [150], which may not always be practically true [151].
@@ 3.2.2 Cross-Attention Fusion
Cross-attention is a relatively more flexible approach to modeling the interactions between data modalities and learning their joint representations. The self-attention layers attend to different modalities at different stages of data processing. Cross-attention allows the model to selectively attend to different modalities based on their relevance to the task [152] and capture complex interactions between the modalities [153].
@@ 3.2.3 Hierarchical Fusion
Hierarchical fusion is a complex approach to combining multiple modalities. For instance, the Depth-supervised Fusion Transformer for Salient Object Detection (DFTR) employs hierarchical feature extraction to improve salient object detection performance by fusing low-level spatial features and high-level semantic features from different scales [154]. Yang et al.[155] introduced a hierarchical approach to fine-grained classification using a fusion Transformer. Furthermore, the Hierarchical Multimodal Transformer (HMT) for video summarization can capture global dependencies and multi-hop relationships among video frames [156].
@@ 3.2.4 Late Fusion
In late fusion, each data modality is processed independently by its own Transformer model, the branch outputs are concatenated and passed through the final classifier. Late fusion allows the model to capture the unique features of each modality while still learning their joint representation. Sun et al. proposed a multi-modal adaptive late fusion Transformer network for estimating the levels of depression [157]. Their model extracts long-term temporal information from audio and visual data independently and then fuses weights at the end to learn a joint representation of data.
Figure 9: Four different strategies of fusing information from various data modalities in multimodal Transformers are presented.
3.3 Transformers for Processing Oncology Datasets
Transformers have been successfully applied to various tasks in oncology, including cancer screening, diagnosis, prognosis, treatment selection, and prediction of clinical variables [29, 158, 159, 110, 143]. For instance, a Transformer-based model was used to predict the presence and grade of breast cancer using a combination of imaging and genomics data [29]. TransMIL [158], a Transformer model, was proposed to process histopathology images using self-attention to learn and classify histopathology slides by overcoming the challenges faced by multi-instance learning (MIL). Recently, a Transformer and convolution parallel network, TransConv [159], was proposed for automatic brain tumor segmentation using MRI data. Transformers and GNNs have also been combined in MML for early-stage NSCLC prognostic prediction using the patient’s clinical and pathological features and by modeling the patient’s physiological network [110]. Similarly, a multimodal co-attention Transformer was proposed for survival prediction using WSIs and genomic sequences [143]. The authors used a co-attention mechanism to learn the interactions between the two data modalities.
4 MML - Challenges and Opportunities